# Benchmark on Drug Target Interaction Modeling from a Structure Perspective

### v1

@@ 5.1 Our Best Combo of Drug and Protein Encoders
Based on our benchmark results, we summarize the insights of protein and drug encoder usages and propose a light yet effective architecture, which could be treated as new strong baseline for the following explorations. Regarding the proteins, we observe that multi-scale CNNs associated with a mixture of model depths can generally learn the effective protein representations [59, 25, 64], which approximates the language model’s accuracy while having lower memory and computation costs.
Regarding the drug molecules, both GNN and Transformer-based methods, such as MRBDTA [55], MolTrans [20] and MGraphDTA [59] prove promising in DTI tasks. This encourages us to leverage information from hybrid perspectives, i.e., implicit structure (via attention in Transformers) and explicit structure learning (via message passing along edges in GNNs).
We are thus motivated to integrate these powerful modules and shed novel insight into the design philosophy of drug-target interaction modeling. Our model combos are illustrated in Fig. 5, where the multi-scale CNNs and hybrid networks of molecular Transformer and GNNs are adopted to learn the representations of proteins and drugs, respectively. As shown by the graph encoder part in Fig. 5, the hybrid networks augment the differential attention matrix in molecular Transformer with inter-atomic distances and graph adjacency matrix [67], which provides the 3D and 2D molecule conformations to further facilitate atom interaction learning.
Specifically, given the projections of molecular input at an attention head, i.e., , the adjacent matrix , and the inter-atomic distances matrix obtained using RDkit, the augmented attention is calculated as:
where is a row-wise softmax function, and and denote scalars weighting the self-attention, distance, and adjacency matrices, respectively. Besides the implicit and explicit structure learning, we integrate the features from drug SMILES. It is notable that simply utilizing the SMILES representation extracted from a transformer for downstream tasks does not perform as well as GNN. To align with the protein embedding paradigm, we adopt a simple CNN to unearth potential SMILES information, as suggested in [68]. Subsequently, due to the fact that cross-attention is more complex and hard to optimize, we implement a straightforward attention mechanism to integrate the representations of the drug graph and SMILES, denoted as and , respectively, using a weighting parameter , as follows:
@@
Benchmark Comparison to State-of-the-Art Frameworks. We compare the proposed combos with the SOTA frameworks in Tables 2 and 3, and Figures 3 and 4. It is observed that our model consistently achieves the best performance in the regression tasks across three datasets and nearly outperforms most methods in classification tasks. By leveraging the physical conformation information from the molecular graph, our combos converge faster than the other two Transformer-based methods, MRBDTA [55] and TDGraphDTA [25], particularly on the KIBA dataset. Moreover, our model uses three times less peak memory and fewer parameters than other Transformer-based methods, enabling faster computation and reduced storage requirements
@@ Appendix A Related Works
GNN-based Methods GNNs play a crucial role in mining the intricate features of drug molecules for drug-target prediction. Numerious models, including Graph Convolutional Network (GCN), Graph Isomorphism Network (GIN), and Graph Attention Network (GAT) have been utilized [18, 35, 38, 69, 69, 61] to process and enhance drug features. Additionally, MGraphDTA [59] employs a multi-scale GNN architecture, while DeepGLSTM [29] leverages parallel GNN structures for drug representation. DeepNC integrates advanced techniques from generalized aggregation networks [70] and hypergraph convolution [71] to improve feature extraction. BACPI [57] develops a bi-directional attention network to integrate the representations of drug molecules and proteins, enhancing their mutual interaction. Besides, BridgeDPI [63] innovates by incorporating bridging nodes between proteins and drugs, utilizing a three-layer GNN for graph embeddings.
Transformer-based Methods Transformers, known for their efficacy in handling sequence data, are extensively applied in drug and protein feature processing. For instance, models like MolTrans [20] and FOTFCPI [41] employ self-attention mechanisms to refine embeddings by focusing on drug and protein substructures. MRBDTA [55] uses multi-head attention and skip connection to enhance drug and protein representation. Additionally, a cross-attention mechanism [32, 42] is employed to facilitate the integration of drug and protein features, enabling effective mutual querying. TDGraphDTA [25] captures contextual relationships between molecular substructures by using a multi-head cross-attention mechanism and graph optimization. Lastly, DrugormerDTI [72] incorporates degree centrality with positional information to highlight the positional relevance of amino acids in proteins.
Input and Featurization Structural information is crucial at the input stage for models such as BridgeDPI [63]. Various libraries, such as DGLGraph [73], DGL-lifeSci [74], and RDKit [46], are employed to process input SMILES of drugs, with RDKit [46] being pivotal for converting SMILE strings into molecular graphs and extracting diverse chemical properties, including chemical bonds, hydrogen presence, electron properties, and so on. Additionally, some approaches [38, 69, 57, 58] incorporate molecular fingerprints [45] to capture local chemical information. For protein sequences, typical pre-processing involves converting amino acid sequences into N-grams [36, 43] or integers [18] sequences. To enhance the expressiveness of embeddings, some models leverage pre-trained Word2Vec [75, 35, 38, 57, 31, 69, 66] or pretrained protein language models [24].
@@ B.1.1 GCN
GraphDTA-GCN [18]: GraphDTA-GCN uses GCN to process the molecular graph, which is derived from SMILES using the RDkit tool, and a simple CNN with integer encoding to handle protein sequences.
GraphCPI-GCN [35]: Similar to GraphDTA, GraphCPI-GCN employs 3-gram encoding with pretrained Word2Vec to process protein sequences, followed by a CNN to handle the protein embeddings.
MGraphDTA [59]: MGraphDTA utilizes a multiscale GCN, inspired by dense connections, and a multiscale CNN to process drug graphs and protein sequences, respectively.
SAGDTA [60]: Similar to GraphDTA, SAGDTA introduces global or hierarchical pooling after GCN to aggregate node representations weightedly.
EmbedDTI [61]: For protein sequences, EmbedDTI leverages GloVe for pretraining amino acid feature embeddings, which are then fed into a CNN. For drugs, it constructs both an atom graph and a substructure graph to capture structural information at different levels, processed by GCN.
DeepGLSTM [29]: DeepGLSTM processes molecular graphs using a parallel GCN module composed of three GCNs with different layers. For protein sequences, it adopts a bi-LSTM.
CPI [31]: CPI processes drug graphs using GCN. The protein sequence is handled via n-gram with integer encoding, followed by a CNN.
DeepNC [30]: DeepNC adopts advanced techniques from generalized aggregation networks and hypergraph convolution, two variants of GCN, to capture the representations of drug. For protein sequences, it uses a simple CNN.
DrugBAN [55]: DrugBAN employs GCN and CNN blocks to encode molecular graph and proteins, respectively. Then they use a bilinear attention network module to learn local interactions between the representations of drugs and proteins.
BridgeDPI [63]: BridgeDPI innovates by constructing a learnable drug–protein association network, which is processed using a three-layer GNN for graph embeddings. The learned representations for drug and protein pairs are then concatenated for further processing.
ColdDTA [64]: ColdDTA removes the subgraphs of drugs. For the model, they adopt the dense GCN and multiscale CNN from MGraphDTA as the encoders for drugs and proteins, respectively. Additionally, an attention-based method is developed to integrate representations for improved prediction.
IMAEN [65]: IMAEN employs a molecular augmentation mechanism to enhance molecular structures by fully aggregating molecular node neighborhood information. It then uses multiscale GCN and CNN for drug and protein processing, respectively.
GanDTI [58]: Inspired by residual networks, GanDTI add the input drug fingerprints to the output of three GCN layers as graph node features and use summation to get the final drug representation.
@@ B.1.2 GAT
GraphDTA-GAT [18]: GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
GraphDTA-GATGCN [18]: GraphDTA-GATGCN adopts a combination of GAT and GCN as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
GraphCPI-GAT [35]: GraphDTA-CPI adopts a GAT as the encoder for drugs, while other components remain the same as in GraphCPI-GCN.
GraphCPI-GATGCN [35]: GraphCPI-GATGCN adopts a combination of GAT and GCN as the encoder for drugs, while other components remain the same as in GraphCPI-GCN.
BACPI [57]: BACPI adopts a GAT and a CNN for the features of the fingerprints and protein sequence, respectively. These features are then fed into a bi-directional attention neural network to obtain integrated representations.
PGraphDTA-CNN [24]: PGraphDTA-CNN is a straightforward method that utilizes GAT for drug feature extraction and CNN for protein sequences.
@@ B.2 GIN
GraphDTA-GIN [18]: GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
GraphCPI-GIN [35]: GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
SubMDTA [36]: SubMDTA utilizes a pretrained GIN encoder obtained through contrastive learning for the molecular graph. For protein sequences, it employs N-gram embedding with different N to extract features at various scales, which are then processed by a BiLSTM.
@@ B.3.1 Self-attention
AMMVF [38]: AWMVF introduces the multi-head mechanism to GAT to learn features in different spaces, and the update function is obtained through the concatenation of different heads’ outputs.
IIFDTI [66]: IIFDTI model attains the drug matrix and protein matrix and inputs them to the bi-directional encoder-decoder block, which considers both the drug and target directions. The decoder is mainly composed of multi-head attention.
MolTrans [20]: MolTrans uses transformer encoder layers to augment the embedding of sub-structure sequences of proteins and drugs.
FOTFCPI [41]: Similar to MolTrans, FOTFCPI uses transformer encoder layers to extract the features of protein and drug fragments after the embedding layers.
TransformerCPI [21]: TransformerCPI uses the decoder module of Transformer, which takes in the atom sequence embedding processed by GCN and the protein sequence embedding processed by word2vec and 1D CNN.
MRBDTA [55]: In MRBDTA, after the embedding layer, drug sequences are directly fed into a block consisting of three Transformer encoders. The first encoder has a linear layer before it and the following two encoders are parallel. The protein sequence is also processed by a block with similar structure.
@@ B.3.2 Cross attention
CSDTI[32]: CSDTI use cross attention to fuse the deep representations of drugs and proteins. Specifically, the different projections of protein feature are used as key and value respectively while the projection of drug feature is used as query.
TDGraphDTA[25]: TDGraphDTA use a multi-head cross-attention mechanism with two attention heads. Both drug and protein features are linearly transformed into query, key and value matrices. One cross attention layer uses a drug query matrix, a protein key matrix, and a protein value matrix, while its parallel counterparts use the rest of the matrices. The outputs of these two layers are concatenated and fed into MLP to get the final output.
@@ Appendix C Datasets Descriptions
In this subsection, we provide a detailed description of the datasets for both the regression task and classification task. The statistical characteristics of the datasets are summarized in Table 4.
@@
| Number of total samples | 25772 | 117657 | 52274 | 6728 | 7786 | 35021 |
@@ Appendix D Evaluation Metrics
We adopt distinct sets of metrics to evaluate the classification and regression tasks. In particular, considering the classification task, we utilize the common metrics including Area Under Receiver Operating Characteristic Curve (ROC-AUC), Precision-Recall Area Under Curve (PR-AUC), LogAUC, accuracy, precision, recall, and F1 score. For the continuous binding affinity regression, we benchmark the models using metrics of mean squared error (MSE), mean absolute error (MAE), coefficient of determination (R2), Pearson correlation coefficient, Concordance Index (CI), and Spearman correlation coefficient. Each of these metrics offers unique insights into different aspects of model performance, allowing us to assess predictive accuracy, correlation with observed values, and consistency in ranking predictions.
Appendix E Greedy Search on Hyperparameter
Here we present the detailed experiment on hyperparameters with full metrics. The detailed review of the hyperparameters associated with the selected models is in Table 5. The detailed greedy hyperparameter seraching results on Davis and Human datasets are in Table 6.

### v2

@@ 5.1 Our Best Combo of Drug and Protein Encoders
Deriving Combo from Benchmark Insights. Based on our benchmark results, we summarize the insights of protein and drug encoder usages and propose a light yet effective architecture, which could be treated as a new strong baseline for future explorations. Regarding the proteins, we observe that multi-scale CNNs associated with a mixture of model depths can generally learn the effective protein representations (yang2022MGraphDTA; zhu2023tdgraph; fang2023colddta), which approximate the language model’s accuracy while having lower memory and computation costs. Regarding the drug molecules, both GNN- and Transformer-based methods, such as MRBDTA (zhang2022predicting), MolTrans (huang2021moltrans), and MGraphDTA (yang2022MGraphDTA) prove promising in DTI tasks. This encourages us to leverage information from hybrid perspectives, i.e., implicit structure (via attention in Transformers) and explicit structure learning (via message passing along edges in GNNs).
Our model design, illustrated in Fig. 5, integrates these components. Specifically, for drug graphs, we adopt a hybrid network that augments the self-attention mechanism with inter-atomic distances and graph adjacency matrices (maziarka2020molecule), incorporating both 2D and 3D molecular structural information. Given the projections of molecular input at an attention head, i.e., the adjacent matrix , and the inter-atomic distances matrix obtained using RDkit, the augmented attention is calculated as follows:
where is a row-wise softmax function, and and denote scalars weighting the self-attention, distance, and adjacency matrices, respectively. Besides the implicit and explicit structure learning, we integrate the features from drug SMILES. It is notable that simply utilizing the SMILES representation extracted from a Transformer for downstream tasks does not perform as well as GNN. To align with the protein embedding paradigm, we adopt a simple CNN to unearth potential SMILES information, as suggested in zhao2022hyperattentiondti. Subsequently, due to the fact that cross-attention is more complex and hard to optimize, we implement a straightforward attention mechanism to integrate the representations of the drug graph and SMILES, denoted as and , respectively, using a weighting parameter , as follows:
@@
Novelty. As opposed to the previous strategy, which heuristically stacked a large amount of modules of different types, our model design is driven by systematic benchmarking and empirical insights. Through extensive experiments under fair and controlled conditions, we identify key encoder and featurization strategies that consistently outperform others. Notably, we disentangle and quantify the distinct molecular features. For instance, atomic properties and hydrogen information significantly enhance predictive performance, while adding electron properties may introduce noise.
Thus, our combo is not merely an ad hoc combination but a carefully validated design that effectively obtains a superior balance between accuracy and computational efficiency on both classification and regression tasks. The clear empirical guidance to model design offered through this study helps to establish a more principled framework for future work in drug-target interaction modeling and provides a robust, reproducible new baseline for the community.
Benchmark Comparison to State-of-the-Art Frameworks. We compare the proposed combos with the SOTA frameworks in Tables˜1 and 2, and Figures˜4 and 4. It is observed that our model consistently achieves the best performance in the regression tasks across three datasets and nearly outperforms most methods in classification tasks. By leveraging the physical conformation information from the molecular graph, our combos converge faster than the other two Transformer-based methods, MRBDTA (zhang2022predicting) and TDGraphDTA (zhu2023tdgraph), particularly on the KIBA dataset. Moreover, our model uses three times less peak memory and fewer parameters than other Transformer-based methods, enabling faster computation and reduced storage requirements.
@@ Appendix A Related Works
GNN-based Methods GNNs play a crucial role in mining the intricate features of drug molecules for drug-target prediction. Numerous models, including Graph Convolutional Network (GCN), Graph Isomorphism Network (GIN), and Graph Attention Network (GAT), have been utilized (nguyen2021graphdta; quan2019graphcpi; wang2023ammvf; lin2020deepgs; jin2021embeddti) to process and enhance drug features. Additionally, MGraphDTA (yang2022MGraphDTA) employs a multi-scale GNN architecture, while DeepGLSTM (mukherjee2022deepglstm) leverages parallel GNN structures for drug representation. DeepNC integrates advanced techniques from generalized aggregation networks (li2020deepergcn) and hypergraph convolution (BAI2021107637) to improve feature extraction. BACPI (li2022bacpi) develops a bi-directional attention network to integrate the representations of drug molecules and proteins, enhancing their mutual interaction. Besides, BridgeDPI (wu2022bridgedpi) innovates by incorporating bridging nodes between proteins and drugs, utilizing a three-layer GNN for graph embeddings.
Transformer-based Methods Transformers, known for their efficacy in handling sequence data, are extensively applied in drug and protein feature processing. For instance, models like MolTrans (huang2021moltrans) and FOTFCPI (yin2024fotf) employ self-attention mechanisms to refine embeddings by focusing on drug and protein substructures. MRBDTA (zhang2022predicting) uses multi-head attention and skip connection to enhance drug and protein representation. Additionally, a cross-attention mechanism (pan2023csdti; kurata2022ican) is employed to facilitate the integration of drug and protein features, enabling effective mutual querying. TDGraphDTA (zhu2023tdgraph) captures contextual relationships between molecular substructures by using a multi-head cross-attention mechanism and graph optimization. Lastly, DrugormerDTI (hu2023drugormerdti) incorporates degree centrality with positional information to highlight the positional relevance of amino acids in proteins.
Input and Featurization Structural information is crucial at the input stage for models such as BridgeDPI (wu2022bridgedpi). Various libraries, such as DGLGraph (wang2019DGL), DGL-lifeSci (li2021dgl-lifesci), and RDKit (landrum2006rdkit), are employed to process input SMILES of drugs, with RDKit (landrum2006rdkit) being pivotal for converting SMILE strings into molecular graphs and extracting diverse chemical properties, including chemical bonds, hydrogen presence, electron properties, and so on. Additionally, some approaches (wang2023ammvf; lin2020deepgs; li2022bacpi; wang2021gandti) incorporate molecular fingerprints (doi:10.1021/ci100050tecfp) to capture local chemical information. For protein sequences, typical preprocessing involves converting amino acid sequences into N-grams (pan2023submdta; 10.1093/bioinformatics/bti801) or integer (nguyen2021graphdta) sequences. To enhance the expressiveness of embeddings, some models leverage pre-trained Word2Vec (10.5555/2999792.2999959word2vec; quan2019graphcpi; wang2023ammvf; li2022bacpi; tsubaki2019compound; lin2020deepgs; cheng2022iifdti) or pre-trained protein language models (bal2024pgraphdta).
@@ B.1.1 GCN
GraphDTA-GCN (nguyen2021graphdta): GraphDTA-GCN uses GCN to process the molecular graph, which is derived from SMILES using the RDkit tool, and a simple CNN with integer encoding to handle protein sequences.
GraphCPI-GCN (quan2019graphcpi): Similar to GraphDTA, GraphCPI-GCN employs 3-gram encoding with pretrained Word2Vec to process protein sequences, followed by a CNN to handle the protein embeddings.
MGraphDTA (yang2022MGraphDTA): MGraphDTA utilizes a multiscale GCN, inspired by dense connections, and a multiscale CNN to process drug graphs and protein sequences, respectively.
SAGDTA (zhang2021sag): Similar to GraphDTA, SAGDTA introduces global or hierarchical pooling after GCN to aggregate node representations weightedly.
EmbedDTI (jin2021embeddti): For protein sequences, EmbedDTI leverages GloVe for pretraining amino acid feature embeddings, which are then fed into a CNN. For drugs, it constructs both an atom graph and a substructure graph to capture structural information at different levels, processed by GCN.
DeepGLSTM (mukherjee2022deepglstm): DeepGLSTM processes molecular graphs using a parallel GCN module composed of three GCNs with different layers. For protein sequences, it adopts a bi-LSTM.
CPI (tsubaki2019compound): CPI processes drug graphs using GCN. The protein sequence is handled via n-gram with integer encoding, followed by a CNN.
DeepNC (tran2022deepnc): DeepNC adopts advanced techniques from generalized aggregation networks and hypergraph convolution, two variants of GCN, to capture the representations of drugs. For protein sequences, it uses a simple CNN.
DrugBAN (zhang2022predicting): DrugBAN employs GCN and CNN blocks to encode molecular graphs and proteins, respectively. Then they use a bilinear attention network module to learn local interactions between the representations of drugs and proteins.
BridgeDPI (wu2022bridgedpi): BridgeDPI innovates by constructing a learnable drug–protein association network, which is processed using a three-layer GNN for graph embeddings. The learned representations for drug and protein pairs are then concatenated for further processing.
ColdDTA (fang2023colddta): ColdDTA removes the subgraphs of drugs. For the model, they adopt the dense GCN and multiscale CNN from MGraphDTA as the encoders for drugs and proteins, respectively. Additionally, an attention-based method is developed to integrate representations for improved prediction.
IMAEN (zhang2024imaen): IMAEN employs a molecular augmentation mechanism to enhance molecular structures by fully aggregating molecular node neighborhood information. It then uses multiscale GCN and CNN for drug and protein processing, respectively.
GanDTI (wang2021gandti): Inspired by residual networks, GanDTI adds the input drug fingerprints to the output of three GCN layers as graph node features and uses summation to get the final drug representation.
@@ B.1.2 GAT
GraphDTA-GAT (nguyen2021graphdta): GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
GraphDTA-GATGCN (nguyen2021graphdta): GraphDTA-GATGCN adopts a combination of GAT and GCN as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
GraphCPI-GAT (quan2019graphcpi): GraphDTA-CPI adopts a GAT as the encoder for drugs, while other components remain the same as in GraphCPI-GCN.
GraphCPI-GATGCN (quan2019graphcpi): GraphCPI-GATGCN adopts a combination of GAT and GCN as the encoder for drugs, while other components remain the same as in GraphCPI-GCN.
BACPI (li2022bacpi): BACPI adopts a GAT and a CNN for the features of the fingerprints and protein sequence, respectively. These features are then fed into a bidirectional attention neural network to obtain integrated representations.
PGraphDTA-CNN (bal2024pgraphdta): PGraphDTA-CNN is a straightforward method that utilizes GAT for drug feature extraction and CNN for protein sequences.
@@ B.2 GIN
GraphDTA-GIN (nguyen2021graphdta): GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
GraphCPI-GIN (quan2019graphcpi): GraphDTA-GAT adopts a GAT as the encoder for drugs, while other components remain the same as in GraphDTA-GCN.
SubMDTA (pan2023submdta): SubMDTA utilizes a pretrained GIN encoder obtained through contrastive learning for the molecular graph. For protein sequences, it employs N-gram embedding with different N to extract features at various scales, which are then processed by a BiLSTM.
@@ B.3.1 Self-attention
AMMVF (wang2023ammvf): AWMVF introduces the multi-head mechanism to GAT to learn features in different spaces, and the update function is obtained through the concatenation of different heads’ outputs.
IIFDTI (cheng2022iifdti): IIFDTI model attains the drug matrix and protein matrix and inputs them to the bi-directional encoder-decoder block, which considers both the drug and target directions. The decoder is mainly composed of multi-head attention.
MolTrans (huang2021moltrans): MolTrans uses Transformer encoder layers to augment the embedding of substructure sequences of proteins and drugs.
FOTFCPI (yin2024fotf): Similar to MolTrans, FOTFCPI uses Transformer encoder layers to extract the features of protein and drug fragments after the embedding layers.
TransformerCPI (chen2020Transformercpi): TransformerCPI uses the decoder module of Transformer, which takes in the atom sequence embedding processed by GCN and the protein sequence embedding processed by word2vec and 1D CNN.
MRBDTA (zhang2022predicting): In MRBDTA, after the embedding layer, drug sequences are directly fed into a block consisting of three Transformer encoders. The first encoder has a linear layer before it and the following two encoders are parallel. The protein sequence is also processed by a block with a similar structure.
@@ B.3.2 Cross attention
CSDTI(pan2023csdti): CSDTI uses cross-attention to fuse the deep representations of drugs and proteins. Specifically, the different projections of protein features are used as keys and values, respectively, while the projection of drug features is used as a query.
TDGraphDTA(zhu2023tdgraph): TDGraphDTA uses a multi-head cross-attention mechanism with two attention heads. Both drug and protein features are linearly transformed into query, key, and value matrices. One cross-attention layer uses a drug query matrix, a protein key matrix, and a protein value matrix, while its parallel counterparts use the rest of the matrices. The outputs of these two layers are concatenated and fed into MLP to get the final output.
@@ Appendix C Datasets Descriptions
In this subsection, we provide a detailed description of the datasets for both the regression task and classification task. The statistical characteristics of the datasets are summarized in Table 3. Here we present the statistics after we cleaned the data as described in Section 3.
@@
| Number of total samples | 25772 | 117657 | 52284 | 5997 | 6552 | 34740 |
@@ Appendix D Evaluation Metrics
We adopt distinct sets of metrics to evaluate the classification and regression tasks. In particular, considering the classification task, we utilize the common metrics, including Area Under Receiver Operating Characteristic Curve (ROC-AUC), Precision-Recall Area Under Curve (PR-AUC), LogAUC, accuracy, precision, recall, and F1 score. For the continuous binding affinity regression, we benchmark the models using metrics of mean squared error (MSE), mean absolute error (MAE), coefficient of determination (R2), Pearson correlation coefficient, concordance index (CI), and Spearman correlation coefficient. Each of these metrics offers unique insights into different aspects of model performance, allowing us to assess predictive accuracy, correlation with observed values, and consistency in ranking predictions.
Appendix E Original Hyperparameter
To have a basic understanding of hyperparameters before greedy search and to find the optimized setting for each model, we summarize the hyperparameters reported in the corresponding paper or codes in Table 4.